OpenStudy (anonymous):

Find the value of 3 over the square root of 3. 3 3 times the square root of 3. the square root of 3. the square root of 3 over 3.

OpenStudy (shamil98):

do you mean?.. \[\frac{ 3 }{ \sqrt{3} }\]

OpenStudy (anonymous):

yeah

OpenStudy (shamil98):

you need to get rid of the square root in the denominator so \[\frac{ 3 }{ \sqrt{3} } \times \frac{ \sqrt{3} }{ \sqrt{3} }\]

OpenStudy (shamil98):

\[\frac{ 3\sqrt{3} }{ 3 }\]

OpenStudy (shamil98):

any number in the square root symbol multiplied byitself is simply that number inside the square roots symbol for example \[\sqrt{4} \times \sqrt{4 } = 4\]

OpenStudy (anonymous):

so its \[\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

??

OpenStudy (shamil98):

It is simply √3

OpenStudy (shamil98):

the 3/3 = 1 cancel each other out .

OpenStudy (anonymous):

oh okay thanks!